OpenStudy (anonymous):

Compute the number of ordered pairs of positive integers (x,y) such that 3x + 5y = 500 Hint: you do not want to list them all, but you want to find a creative way to count how many first-quadrant ordered pairs satisfy this equation.

OpenStudy (anonymous):

i know it says creative way but...

OpenStudy (anonymous):

how do I find them ALLL??

OpenStudy (anonymous):

is it like, find the x and y intercepts and then the slope and then just find your way from there??

jimthompson5910 (jim_thompson5910):

what happens when x = 0 ?

OpenStudy (anonymous):

y = 100

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

when y is 0, x is ...

jimthompson5910 (jim_thompson5910):

so (0,100) is one solution BUT x is not positive here

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

what is the slope of this line?

OpenStudy (anonymous):

um 5y= -3x+500 y= -3/5x + 100

OpenStudy (anonymous):

so slope is -3/5x

jimthompson5910 (jim_thompson5910):

just -3/5 it's the number out front of the x

OpenStudy (anonymous):

oh wait, yeah, oops

jimthompson5910 (jim_thompson5910):

slope = -3/5 = rise/run = (change in y)/(change in x) what this means is that when x increases by 5, y decreases by 3 so (0,100) would turn into (5,97) following the rule above notice how 0 turned into 5 (x increased by 5) notice how 100 turned into 97 (y decreased by 3)

OpenStudy (anonymous):

mmmhmmm

OpenStudy (anonymous):

what about..

jimthompson5910 (jim_thompson5910):

the question is: when does y become negative?

OpenStudy (anonymous):

when... 33 ordered pairs later..?

OpenStudy (anonymous):

but during that, x becomes negative

jimthompson5910 (jim_thompson5910):

more like 34 100 - 34*3 = 100 - 102 = -2

jimthompson5910 (jim_thompson5910):

oh I see what you mean

OpenStudy (anonymous):

oh, yes

OpenStudy (anonymous):

so would so since x becomes negative faster than x, do i do something like 100-5x or something?

OpenStudy (anonymous):

agh wait, i meant

OpenStudy (anonymous):

so 100- 20x5...?

OpenStudy (anonymous):

and then taking out the intercepts would mean I have 18 ordered pairs

jimthompson5910 (jim_thompson5910):

to generate any ordered pair (x,y) where you want this point to be in Q1, use this rule x = 0 + 5*k y = 100 - 3*k k is an integer k ranges from k = 1 to k = 33 if k < 1 or k > 33, then (x,y) will not be in Q1 (it will be on an axis or in another quadrant)

jimthompson5910 (jim_thompson5910):

so this shows that there are 33 ordered pairs (x,y) where x and y are positive integers

OpenStudy (anonymous):

ohwait but what about the x being negative?

jimthompson5910 (jim_thompson5910):

that happens when k < 1 which I mentioned above

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

thank you (like usual)

jimthompson5910 (jim_thompson5910):

np